# Molecular De Novo Design through Transformer-based Reinforcement Learning

### v1

@@ Introduction
The vast expanse of chemical space, encompassing an order of magnitude from $10^{60}-10^{100}$ possible synthetically feasible molecules (Engelmore & Morgan, 1986), presents formidable obstacles to drug discovery endeavors. In this colossal landscape, the task of pinpointing a molecule that simultaneously meets the prerequisites for bioactivity, drug metabolism and pharmacokinetic (DMPK) profile, and synthetic accessibility becomes an undertaking similar to the proverbial search for a needle in a haystack. Pioneering de novo design algorithms (Clancey, 1983) have attempted to address this by employing virtual strategies to design and evaluate molecules, thereby condensing the vast chemical space into a more navigable realm for exploration.
Traditional de novo design models, based on Recurrent Neural Networks (RNNs), have proven effective in molecule generation tasks (Clancey, 1984; Robinson, 1980a). However, RNNs possess inherent architectural limitations, notably in their capability to capture long-term dependencies in sequential data, which can be particularly detrimental when modeling complex molecular structures. Recently, the Transformer architecture has emerged as a powerful alternative to RNNs in sequence modeling tasks across various domains. Some of the key advantages of Transformers over RNNs include:
@@
Drawing inspiration from previous work that employed RNNs and reinforcement learning for molecular optimization (Clancey, 1983), our approach distinguishes itself by the adoption and fine-tuning of the Transformer architecture, ensuring superior handling of long-sequence data and paving the way for innovative breakthroughs in the realm of molecular design.
@@
Early de novo design algorithms were structure-based, aiming to grow ligands to fit the binding pocket of the target (Robinson, 1980b; Hasling et al., 1984). However, these methods often generated molecules with poor DMPK properties and could be synthetically intractable. Ligand-based approaches were introduced to create a vast virtual library of chemical structures and then searched with a scoring function (Hasling et al., 1983; Rice, 1986).
Recently, generative models such as RNN-based methods have been used for de novo design of molecules (Clancey, 1979; 2021). They have shown success in tasks like learning the underlying probability distribution over a large set of chemical structures, reducing the search over chemical space to only molecules seen as reasonable. Further fine-tuning of the models was done using reinforcement learning (RL) (Bouville, 2008), which showed considerable improvement over the initial model.
Despite these advancements, challenges such as capturing long-term dependencies in the sequence data persist. The Transformer architecture (NASA, 2015), known for its self-attention mechanism and ability to handle long sequences, has been highly successful in several sequence prediction tasks across domains. Motivated by these successes, we propose the use of Transformer-based architectures in place of RNNs for molecular de novo design.
@@
Transformers Overview Transformers are a neural network architecture designed to process sequential data, while also accounting for the importance of each input in relation to the others, despite their position in the sequence (Goodfellow et al., 2016). They manage to do this by the introduction of an attention mechanism that assesses the significance of each input in the sequence (Figure 1). At any given step $t$, the transformer state at $t$ is influenced by all previous inputs $x^{1},\ldots,x^{t-1}$ and the current input $x^{t}$. The transformer’s ability to selectively focus on the parts of the input sequence that are most relevant for each step makes them especially well suited for tasks in the field of natural language processing. Sequences of words can be encoded into one-hot vectors with a length equivalent to our vocabulary size $X$. We may add two extra tokens, GO and EOS, to signify the beginning and end of a sequence, respectively.
Learning the data Training a Transformer for sequence modeling typically involves using maximum likelihood estimation to predict the next token $x^{t}$ in the target sequence, given tokens from the previous steps (Figure 1). The model generates a probability distribution at every step, representing the likely next character, and the objective is to maximize the likelihood assigned to the correct token:
@@
Sample Efficiency
The primary metric, AUC top-K, emphasizes not just optimization capability but also sample efficiency. High values in this metric would imply fewer oracle calls, thereby maximizing performance with limited data evaluations. REINVENT-Trans, achieving high scores in several oracles, suggests its effectiveness in rapidly optimizing toward desirable molecular properties without exhaustive database searches.
Future Implications
These results showcase the potential of transformers in the domain of molecular generation. Given the success of transformer models in natural language processing tasks, it’s no surprise that their capabilities translate effectively into molecular representations as well.
However, it’s essential to consider the versatility of the tasks presented. Some models might specialize in particular tasks but may not be universally applicable. Hence, it’s beneficial to have an ensemble or a selection mechanism based on the specific task at hand.

### v3

@@ Introduction
The vast expanse of chemical space, encompassing an order of magnitude from $10^{60}-10^{100}$ possible synthetically feasible molecules (Schneider & Fechner, 2005), presents formidable obstacles to drug discovery endeavors. In this colossal landscape, the task of pinpointing a molecule that simultaneously meets the prerequisites for bioactivity, drug metabolism and pharmacokinetic (DMPK) profile, and synthetic accessibility becomes an undertaking similar to the proverbial search for a needle in a haystack. Pioneering de novo design algorithms (Böhm, 1992; Gillet et al., 1994) have attempted to address this by employing virtual strategies to design and evaluate molecules, thereby condensing the vast chemical space into a more navigable realm for exploration.
Traditional de novo design models, based on Recurrent Neural Networks (RNNs), have proven effective in molecule generation tasks (Olivecrona et al., 2017). However, RNNs possess inherent architectural limitations, notably in their capability to capture long-term dependencies in sequential data, which can be particularly detrimental when modeling complex molecular structures. Recently, the Transformer architecture has emerged as a powerful alternative to RNNs in sequence modeling tasks across various domains. Some of the key advantages of Transformers over RNNs include:
@@
Drawing inspiration from previous work that employed RNNs and reinforcement learning for molecular optimization (Olivecrona et al., 2017), our approach distinguishes itself by the adoption and fine-tuning of the Transformer architecture, ensuring superior handling of long-sequence data and paving the way for innovative breakthroughs in the realm of molecular design.
@@
Early de novo design algorithms were structure-based, aiming to grow ligands to fit the binding pocket of the target (Böhm, 1992; Gillet et al., 1994). However, these methods often generated molecules with poor DMPK properties and could be synthetically intractable. Ligand-based approaches were introduced to create a vast virtual library of chemical structures and then searched with a scoring function (Ruddigkeit et al., 2013; Hartenfeller et al., 2012).
Recently, generative models such as RNN-based methods have been used for de novo design of molecules (Segler et al., 2017; Gómez-Bombarelli et al., 2018; Yu et al., 2017). They have shown success in tasks like learning the underlying probability distribution over a large set of chemical structures, reducing the search over chemical space to only molecules seen as reasonable. Further fine-tuning of the models was done using reinforcement learning (RL) (Jaques et al., 2017), which showed considerable improvement over the initial model.
Despite these advancements, challenges such as capturing long-term dependencies in the sequence data persist. The Transformer architecture (Vaswani et al., 2023), known for its self-attention mechanism and ability to handle long sequences, has been highly successful in several sequence prediction tasks across domains. Motivated by these successes, we propose the use of Transformer-based architectures in place of RNNs for molecular de novo design.
@@
Transformers Overview Transformers are a neural network architecture designed to process sequential data, while also accounting for the importance of each input in relation to the others, despite their position in the sequence (Vaswani et al., 2023). They manage to do this by the introduction of an attention mechanism that assesses the significance of each input in the sequence (Figure 1). At any given step $t$, the transformer state at $t$ is influenced by all previous inputs $x^{1},\ldots,x^{t-1}$ and the current input $x^{t}$. The transformer’s ability to selectively focus on the parts of the input sequence that are most relevant for each step makes them especially well suited for tasks in the field of natural language processing. Sequences of words can be encoded into one-hot vectors with a length equivalent to our vocabulary size $X$. We may add two extra tokens, GO and EOS, to signify the beginning and end of a sequence, respectively.
Learning to model the data Training a Transformer for sequence modeling typically involves using maximum likelihood estimation to predict the next token $x^{t}$ in the target sequence, given tokens from the previous steps (Figure 1). The model generates a probability distribution at every step, representing the likely next character, and the objective is to maximize the likelihood assigned to the correct token:
@@
In order to better investigate in the ability of our method in long sequence generation, we did the following ablation study.